# spacing=nonfrench

### v2

@@ II.4 Coupling to a secondary reaction
Thus, energy can be directly extracted from a temperature gradient using these reaction-diffusion systems. However, the design of a coupled chemical reaction network implies the ability to propagate energy input to the secondary processes[3, 14, 15]. We further investigate the coupling of the minimalist network in Eq. (1) to maintain nonequilibrium secondary processes.
@@ III Conclusions
The entropy dissipated by the reaction-diffusion system is distributed unequally among each process. A net difference can be observed between the diffusion-dominated and reaction-dominated systems. The desired regime can be tailored by scaling the system to a critical system length $\mathcal{L}_{0}$. It should be noted that the traditional “local equilibrium” approximation reduces to the limit case of fast reactions, that is, to large systems whose physical length is well above $\mathcal{L}_{0}$. This regime is suboptimal for generating nonequilibrium reaction fluxes; it is the worst case for extracting chemical energy from the temperature gradient, as it leads to minimization of the chemical flux.
@@ Appendix A Numerical tools
where $\bar{x}_{0}$ is the position at which $\varphi=0$ (see D.2). In all simulations, $\omega=10^{-6}$ was used. When necessary, the total simulation time was increased until the condition of Eq. (25z) was met.
The calculations were performed using XMDS2[45]. In each simulation, a steady temperature gradient was imposed with either a Heaviside, linear, or mixed profile. In all models, the influence of temperature on the chemical reaction was introduced via the Arrhenius relationship for the reaction rate constants. Integration was performed using the 8th (embedded 9th) order adaptive Runge-Kutta method. Spatial derivation was performed on the basis of a discrete cosine transform for the implementation of zero Neumann boundaries, thus describing a closed system.
Appendix B Parameter description
The chemical system is reduced to the following single chemical reaction:
@@ Appendix B Parameter description
with
$K=\frac{k_{+}}{k_{-}}$ (25ab)
Using thermodynamics and Arrhenius relationships, the variation in the constants with temperature can be expressed as:
$K$ $=K^{\infty}e^{-\frac{\Delta_{r}H_{0}}{RT}}$ (25aca)
$k_{+}$ $=k_{+}^{\infty}e^{-\frac{E_{a,+}}{RT}}$ (25acb)
$k_{-}$ $=k_{-}^{\infty}e^{-\frac{E_{a,-}}{RT}}$ (25acc)
with
$\Delta_{r}H_{0}$ $=E_{a,+}-E_{a,-}$ (25acd)
These parameters can be described from a reference temperature $T_{0}$, so that $K_{0}=K(T_{0})$, $k_{+,0}=k_{+}(T_{0})$, $k_{-,0}=k_{-}(T_{0})$, leading to:
$K$ $=K_{0}e^{\frac{\Delta_{r}H_{0}}{RT_{0}}\left(\frac{T-T_{0}}{T}\right)}$ (25ada)
$k_{+}$ $=k_{+,0}e^{\frac{E_{a,+}}{RT_{0}}\left(\frac{T-T_{0}}{T}\right)}$ (25adb)
$k_{-}$ $=k_{-,0}e^{\frac{E_{a,-}}{RT_{0}}\left(\frac{T-T_{0}}{T}\right)}$ (25adc)
This is then nondimensionalized as:
$\bar{k}_{+}$ $=\kappa\rho e^{\beta(1-\gamma)\frac{\theta}{\theta+1}}$ (25aea)
$\bar{k}_{-}$ $=\kappa^{-1}\rho e^{\beta(1+\gamma)\frac{\theta}{\theta+1}}$ (25aeb)
$K$ $=\kappa^{2}e^{2\beta\gamma\frac{\theta}{\theta+1}}$ (25aec)
$\eta$ $=2\beta\gamma$ (25aed)
$\varepsilon_{+}$ $=\beta(1-\gamma)$ (25aee)
$\varepsilon_{-}$ $=\beta(1+\gamma)$ (25aef)
$\beta$ corresponds to the average activation energy, and $\gamma$ corresponds to the energetic asymmetry of the system (see Fig. S5 in SI).
In the two-compartments model, diffusion is modeled as a first-order exchange reaction between two compartments as a function of the geometric mean of the exchange kinetic constant $\delta$:
$\delta_{1}$ $=\lambda\delta$ (25afa)
$\delta_{2}$ $=\lambda^{-1}\delta$ (25afb)
In unstirred systems, the free diffusion of the chemical reactants is fully considered. Neglecting the variation in the diffusion constant with temperature, chemical diffusion is described as a function of the geometric mean of the diffusion constant $d$, and the exchange ratio $\lambda$:
$d_{1}$ $=\lambda d$ (25aga)
$d_{2}$ $=\lambda^{-1}d$ (25agb)
@@ Appendix C Mathematical derivation in the stirred compartments model
The temperature is set to $-\theta_{g}$ in compartment $A$, and to $+\theta_{g}$ in compartment $B$. The initial concentration is $1$ in each compartment, i.e.:
$u_{1a}(0)+u_{2a}(0)$ $=u_{1b}(0)+u_{2b}(0)=1$ (25ah)
The system is described by the following chemical processes:
The dynamics of this system is described by the following set of ODEs:
$\frac{du_{1a}}{d\bar{t}}$ $=\lambda^{-1}\delta\left(-u_{1a}+u_{1b}\right)-\bar{k}_{+,a}u_{1a}+\bar{k}_{-,a}u_{2a}$ (25aiaja)
$=-\varphi_{d1}-\varphi_{ra}$ (25aiajb)
$\frac{du_{1b}}{d\bar{t}}$ $=\lambda^{-1}\delta\left(u_{1a}-u_{1b}\right)-\bar{k}_{+,b}u_{1b}+\bar{k}_{-,b}u_{2b}$ (25aiajc)
$=+\varphi_{d1}-\varphi_{rb}$ (25aiajd)
$\frac{du_{2a}}{d\bar{t}}$ $=\lambda\delta\left(-u_{2a}+u_{2b}\right)+\bar{k}_{+,a}u_{1a}-\bar{k}_{-,a}u_{2a}$ (25aiaje)
$=-\varphi_{d2}+\varphi_{ra}$ (25aiajf)
$\frac{du_{2b}}{d\bar{t}}$ $=\lambda\delta\left(u_{2a}-u_{2b}\right)+\bar{k}_{+,b}u_{1b}-\bar{k}_{-,b}u_{2b}$ (25aiajg)
$=+\varphi_{d2}+\varphi_{rb}$ (25aiajh)
@@
The steady state $u_{ij}^{\infty}$ (with $i\in[1,2]$ and $j\in[a,b]$) can be obtained directly from the resolution of the set of equations (25aiaj), for $\dfrac{du_{ij}}{d\bar{t}}=0$. Each chemical force and flux of any reaction or exchange process $k$ can be expressed as:
$\varphi_{k}$ $=r_{k}^{+}-r_{k}^{-}$ (25aiaka)
$\alpha_{k}$ $=\ln\frac{r_{k}^{+}}{r_{k}^{-}}$ (25aiakb)
@@
$\varphi_{ra}$ $=\bar{k}_{+,a}u^{\infty}_{1a}-\bar{k}_{-,a}u^{\infty}_{2a}$ (25aiala)
$\alpha_{ra}$ $=\ln\left(K_{a}\frac{u^{\infty}_{1a}}{u^{\infty}_{2a}}\right)$ (25aialb)
@@
$\varphi_{rb}$ $=\bar{k}_{+,b}u^{\infty}_{1b}-\bar{k}_{-,b}u^{\infty}_{2b}$ (25aialc)
$\alpha_{rb}$ $=\ln\left(K_{b}\frac{u^{\infty}_{1b}}{u^{\infty}_{2b}}\right)$ (25aiald)
@@
$\varphi_{d1}$ $=\lambda^{-1}\delta(u^{\infty}_{1a}-u^{\infty}_{1b})$ (25aiale)
$\alpha_{d1}$ $=\ln\left(\frac{u^{\infty}_{1a}}{u^{\infty}_{1b}}\right)$ (25aialf)
@@
$\varphi_{d2}$ $=\lambda\delta(u^{\infty}_{2a}-u^{\infty}_{2b})$ (25aialg)
$\alpha_{d2}$ $=\ln\left(\frac{u^{\infty}_{2a}}{u^{\infty}_{2b}}\right)$ (25aialh)
@@
At the steady state, all the chemical fluxes compensate for each other, following Eqs. (25aiaj) such that:
$\varphi$ $=-\varphi_{ra}=\varphi_{rb}=\varphi_{d1}=-\varphi_{d2}$ (25aiam)
@@
$R_{k}$ $=\frac{\alpha_{k}}{\varphi_{k}}$ (25aian)
Each process dissipate entropy as:
$\sigma_{k}$ $=\alpha_{k}\varphi_{k}$ (25aiaoa)
$=R_{k}\varphi_{k}^{2}=R_{k}\varphi^{2}$ (25aiaob)
@@
The full entropy dissipation is then:
$\sigma$ $=\sum_{k}\sigma_{k}$ (25aiapa)
$=\sum_{k}R_{k}\varphi^{2}$ (25aiapb)
This yields the total dissipation of entropy in the system as:
$\sigma_{ra}$ $=\alpha_{ra}\varphi_{ra}=-\alpha_{ra}\varphi$ (25aiaqa)
$=R_{ra}\varphi^{2}$ (25aiaqb)
$\sigma_{rb}$ $=\alpha_{rb}\varphi_{rb}=+\alpha_{rb}\varphi$ (25aiaqc)
$=R_{rb}\varphi^{b}$ (25aiaqd)
$\sigma_{d1}$ $=\alpha_{d1}\varphi_{d1}=+\alpha_{d1}\varphi$ (25aiaqe)
$=R_{d1}\varphi^{2}$ (25aiaqf)
$\sigma_{d2}$ $=\alpha_{d2}\varphi_{d2}=-\alpha_{d2}\varphi$ (25aiaqg)
$=R_{d2}\varphi^{2}$ (25aiaqh)
$\sigma_{r}$ $=\sigma_{ra}+\sigma_{rb}=(\alpha_{rb}-\alpha_{ra})\varphi$ (25aiaqi)
$=R_{r}\varphi^{2}$ (25aiaqj)
$\sigma_{d}$ $=\sigma_{d1}+\sigma_{d2}=(\alpha_{d1}-\alpha_{d2})\varphi$ (25aiaqk)
$=R_{d}\varphi^{2}$ (25aiaql)
$\sigma$ $=\sigma_{r}+\sigma_{d}=(\alpha_{r2}-\alpha_{r1}+\alpha_{d1}-\alpha_{d2})\varphi$ (25aiaqm)
$=\alpha\varphi=R_{\text{tot}}\varphi^{2}$ (25aiaqn)
@@
$R_{\text{tot}}$ $=R_{r}+R_{d}$ (25aiaqo)
$R_{r}$ $=R_{ra}+R_{rb}$ (25aiaqp)
$R_{d}$ $=R_{d1}+R_{d2}$ (25aiaqq)
According to Eqs. (25aial), (25aiam) and (25aec), the total chemical force $\alpha$ can be expressed as:
$\alpha$ $=-\alpha_{ra}+\alpha_{rb}+\alpha_{d1}-\alpha_{d2}$ (25aiara)
$=\ln\left(K_{a}^{-1}\frac{u^{\infty}_{2a}}{u^{\infty}_{1a}}\cdot K_{b}\frac{u^{\infty}_{1b}}{u^{\infty}_{2b}}\cdot\frac{u^{\infty}_{1a}}{u^{\infty}_{1b}}\cdot\frac{u^{\infty}_{2b}}{u^{\infty}_{2a}}\right)$ (25aiarb)
$=\ln\left(\frac{K_{b}}{K_{a}}\right)$ (25aiarc)
$=\left(\frac{\theta_{g}}{\theta_{g}+1}-\frac{\theta_{g}}{\theta_{g}-1}\right)\eta$ (25aiard)
$=\frac{2\eta\theta_{g}}{1-\theta_{g}^{2}}$ (25aiare)
@@
$\alpha$ $=2\eta\theta_{g}+O(\theta_{g}^{3})\approx 2\eta\theta_{g}$ (25aias)
@@ C.2 Expression of the steady chemical flux
$\varphi$ $=\frac{\rho\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\frac{\kappa^{\prime}\eta\theta_{g}}{2}$ (25aiata)
@@
$\kappa^{\prime}$ $=\frac{2\kappa}{1+\kappa^{2}}$ (25aiatb)
@@
$\lambda^{\prime}$ $=\frac{2\kappa\lambda}{1+\kappa^{2}\lambda^{2}}$ (25aiatc)
$\kappa^{\prime}$ and $\lambda^{\prime}$ are parameters that characterize the asymmetry of the chemical reaction and chemical exchange, respectively. The chemical flux in Eq. (25aiata) indicates that the behavior of the system depends on the relative values of $\rho$ and $\lambda^{\prime}\delta$, with the intensity proportional to the reaction asymmetry $\kappa^{\prime}$.
@@
The total dissipation is distributed between the chemical reactions and exchanges:
$\alpha_{r}$ $=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\alpha$ (25aiaua)
$\alpha_{d}$ $=\frac{\rho}{\lambda^{\prime}\delta+\rho}\alpha$ (25aiaub)
@@
$\sigma_{r}$ $=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\sigma$ (25aiava)
$\sigma_{d}$ $=\frac{\rho}{\lambda^{\prime}\delta+\rho}\sigma$ (25aiavb)
Distribution of the chemical reaction dissipation in each compartment
@@
$\alpha_{ra}$ $=-\frac{1+\varepsilon}{2}\alpha_{r}$ (25aiawa)
$\alpha_{rb}$ $=+\frac{1-\varepsilon}{2}\alpha_{r}$ (25aiawb)
@@
$\varepsilon$ $=\beta+\frac{\eta\rho}{2(\lambda^{\prime}\delta+\rho)}\frac{1-\kappa^{2}\lambda^{2}}{1+\kappa^{2}\lambda^{2}}$ (25aiawc)
implying the following distribution of entropy dissipation:
$\sigma_{ra}$ $=\frac{1+\varepsilon\theta_{g}}{2}\sigma_{r}$ (25aiaxa)
$\sigma_{rb}$ $=\frac{1-\varepsilon\theta_{g}}{2}\sigma_{r}$ (25aiaxb)
@@
Distribution of the exchange dissipation for each compound
@@
$\alpha_{d1}$ $=+\frac{\lambda^{2}\kappa^{2}}{1+\lambda^{2}\kappa^{2}}\alpha_{d}$ (25aiaya)
$\alpha_{d2}$ $=-\frac{1}{1+\lambda^{2}\kappa^{2}}\alpha_{d}$ (25aiayb)
implying the following distribution of entropy dissipation:
$\sigma_{d1}$ $=\frac{\lambda^{2}\kappa^{2}}{1+\lambda^{2}\kappa^{2}}\sigma_{d}$ (25aiaza)
$\sigma_{d2}$ $=\frac{1}{1+\lambda^{2}\kappa^{2}}\sigma_{d}$ (25aiazb)

### v3

@@ II.4 Coupling to a secondary reaction
[…]
@@ III Conclusions
From the model established in this work, it is possible to identify the best experimental conditions (i.e in terms of the values of the minimal, mean and maximal temperatures, the shape of the temperature gradient, and the system dimension) for sustaining efficiently a given a chemical system in a nonequilibrium steady state (characterized by a maximal values of entropy production by the targeted chemical process). Moreover, the entropy produced by the reaction-diffusion system is distributed unequally among each process. A net difference can be observed between the diffusion-dominated and reaction-dominated systems. The desired regime can be tailored by scaling the system to a critical system length $\mathcal{L}_{0}$. It should be noted that the traditional “local equilibrium” approximation reduces to the limit case of fast reactions, that is, to large systems whose physical length is well above $\mathcal{L}_{0}$. This regime is suboptimal for generating nonequilibrium reaction fluxes; it is the worst case for extracting chemical energy from the temperature gradient, as it leads to minimization of the chemical flux.
@@ Appendix A Numerical tools
where $\bar{x}_{0}$ is the position at which $\varphi=0$ (see D.2). In all simulations, $\omega=10^{-6}$ was used. When necessary, the total simulation time was increased until the condition of Eq. (4bvbwbx) was met.
The calculations were performed using XMDS2, an open source software for solving PDEs using spectral methods for computing spatial derivatives [47]. In each simulation, a steady temperature gradient was imposed with either a Heaviside, linear, or mixed profile. In all models, the influence of temperature on the chemical reaction was introduced via the Arrhenius relationship for the reaction rate constants. Integration was performed using the 8th (embedded 9th) order adaptive Runge-Kutta method. Spatial derivation was performed on the basis of a discrete cosine transform for the implementation of zero Neumann boundaries, thus describing a closed system.
Appendix B Entropy production due to thermal diffusion
The system is submitted to a temperature gradient. It is thus subjected to a steady heat flux, generating a production of entropy $\dot{S}_{\mathrm{th}}$, nondimensionalized as $\sigma_{\mathrm{th}}$. Assuming a linear gradient from $T_{0}(1-\theta_{g})$ to $T_{0}(1+\theta_{g})$ over a length $l_{g}$, inside a system of length $l_{0}$, this term can be expressed at each position $x$ as [36]:
$T$ $=T_{0}\cdot\left(1+\left(\frac{2\theta_{g}}{l_{g}}x-\theta_{g}\right)\right)$ (4bvbwbya)
$\dot{S}_{\mathrm{th}}(x)$ $=k_{\mathrm{th}}\left(\frac{1}{T}\cdot\frac{\partial T}{\partial x}\right)^{2}$ (4bvbwbyb)
$=k_{\mathrm{th}}\frac{4\theta_{g}^{2}}{(l_{g}(1-\theta_{g})+2\theta_{g}x)^{2}}$ (4bvbwbyc)
$k_{\mathrm{th}}$ being the thermal conductivity of the medium (in J.s-1.m-1.K-1).
On average, the entropy production in the full system is thus:
@@ Appendix B Entropy production due to thermal diffusion
$=\frac{4k_{\mathrm{th}}}{l_{0}l_{g}}\cdot\frac{\theta_{g}^{2}}{1-\theta_{g}^{2}}$ (4bvbwbzb)
This can be nondimensionalized as:
$\sigma_{\mathrm{th}}$ $=\dot{S}_{\mathrm{th}}\cdot\frac{l_{0}^{2}}{Rc_{0}D_{0}}$ (4bvbwcaa)
$=\frac{4\bar{k}_{\mathrm{th}}}{\bar{l}_{g}}\cdot\frac{\theta_{g}^{2}}{1-\theta_{g}^{2}}$ (4bvbwcab)
$\approx\frac{4\bar{k}_{\mathrm{th}}}{\bar{l}_{g}}\theta_{g}^{2}\quad\textrm{for $\theta_{g}\ll 1$}$ (4bvbwcac)
$\textrm{with}\qquad\bar{k}_{\mathrm{th}}$ $=\frac{k_{\mathrm{th}}}{Rc_{0}D_{0}}$ (4bvbwcad)
$\bar{k}_{\mathrm{th}}$ is the nondimensionalized thermal conductivity. For example, in the case of water as the system solvent, $k_{\mathrm{th}}=0.61$ J.s-1.K-1 at $300$ K [48]; with a characteristic reactant concentration $c_{0}=10^{3}$ mol.m-3 and chemical diffusion $D_{0}=10^{-9}$ m2.s-1, $\bar{k}_{\mathrm{th}}=7.3\times 10^{4}$. This value can be compared with $\eta^{2}$, in order to compare the intensity of the thermal entropy production with the chemical entropy production (Eq. (4bvbwcvi). An example system, with $\eta^{2}=1600$ for a reaction enthalpy $\Delta_{r}H=100$ kJ.mol-1 at $300$ K, would imply a chemical entropy production that represents at most 1% of the thermal entropy production.
$\bar{l}_{g}$ represents the fraction on the system where the gradient is established ($1$ for the linear gradient, $\nicefrac{{1}}{{2}}$ for the mixed gradient, and 00 for the Heaviside gradient and a two-compartment system). It shall be noted that the entropy production thus tends to infinity in the case of the Heaviside gradient; this is due to the non-physicality of sustaining an infinitely sharp temperature gradient.
@@ C.1 Steady state
The steady state $u_{ij}^{\infty}$ (with $i\in[1,2]$ and $j\in[a,b]$) can be obtained directly from the resolution of the set of equations (4af), for $\dfrac{du_{ij}}{d\bar{t}}=0$. Each chemical force and flux of any reaction or exchange process $k$ can be expressed as:
$\varphi_{k}$ $=r_{k}^{+}-r_{k}^{-}$ (4bvbwcba)
$\alpha_{k}$ $=\ln\frac{r_{k}^{+}}{r_{k}^{-}}$ (4bvbwcbb)
@@
$\varphi_{ra}$ $=\bar{k}_{+,a}u^{\infty}_{1a}-\bar{k}_{-,a}u^{\infty}_{2a}$ (4bvbwcca)
$\alpha_{ra}$ $=\ln\left(K_{a}\frac{u^{\infty}_{1a}}{u^{\infty}_{2a}}\right)$ (4bvbwccb)
@@
$\varphi_{rb}$ $=\bar{k}_{+,b}u^{\infty}_{1b}-\bar{k}_{-,b}u^{\infty}_{2b}$ (4bvbwccc)
$\alpha_{rb}$ $=\ln\left(K_{b}\frac{u^{\infty}_{1b}}{u^{\infty}_{2b}}\right)$ (4bvbwccd)
@@
$\varphi_{d1}$ $=\lambda^{-1}\delta(u^{\infty}_{1a}-u^{\infty}_{1b})$ (4bvbwcce)
$\alpha_{d1}$ $=\ln\left(\frac{u^{\infty}_{1a}}{u^{\infty}_{1b}}\right)$ (4bvbwccf)
@@
$\varphi_{d2}$ $=\lambda\delta(u^{\infty}_{2a}-u^{\infty}_{2b})$ (4bvbwccg)
$\alpha_{d2}$ $=\ln\left(\frac{u^{\infty}_{2a}}{u^{\infty}_{2b}}\right)$ (4bvbwcch)
@@
At the steady state, all the chemical fluxes compensate for each other, following Eqs. (4af) such that:
$\varphi$ $=-\varphi_{ra}=\varphi_{rb}=\varphi_{d1}=-\varphi_{d2}$ (4bvbwcd)
@@
$R_{k}$ $=\frac{\alpha_{k}}{\varphi_{k}}$ (4bvbwce)
Each process produces entropy as:
$\sigma_{k}$ $=\alpha_{k}\varphi_{k}$ (4bvbwcfa)
$=R_{k}\varphi_{k}^{2}=R_{k}\varphi^{2}$ (4bvbwcfb)
@@
The full entropy production is then:
$\sigma$ $=\sum_{k}\sigma_{k}$ (4bvbwcga)
$=\sum_{k}R_{k}\varphi^{2}$ (4bvbwcgb)
This yields the total production of entropy in the system as:
$\sigma_{ra}$ $=\alpha_{ra}\varphi_{ra}=-\alpha_{ra}\varphi$ (4bvbwcha)
$=R_{ra}\varphi^{2}$ (4bvbwchb)
$\sigma_{rb}$ $=\alpha_{rb}\varphi_{rb}=+\alpha_{rb}\varphi$ (4bvbwchc)
$=R_{rb}\varphi^{b}$ (4bvbwchd)
$\sigma_{d1}$ $=\alpha_{d1}\varphi_{d1}=+\alpha_{d1}\varphi$ (4bvbwche)
$=R_{d1}\varphi^{2}$ (4bvbwchf)
$\sigma_{d2}$ $=\alpha_{d2}\varphi_{d2}=-\alpha_{d2}\varphi$ (4bvbwchg)
$=R_{d2}\varphi^{2}$ (4bvbwchh)
$\sigma_{r}$ $=\sigma_{ra}+\sigma_{rb}=(\alpha_{rb}-\alpha_{ra})\varphi$ (4bvbwchi)
$=R_{r}\varphi^{2}$ (4bvbwchj)
$\sigma_{d}$ $=\sigma_{d1}+\sigma_{d2}=(\alpha_{d1}-\alpha_{d2})\varphi$ (4bvbwchk)
$=R_{d}\varphi^{2}$ (4bvbwchl)
$\sigma$ $=\sigma_{r}+\sigma_{d}=(\alpha_{r2}-\alpha_{r1}+\alpha_{d1}-\alpha_{d2})\varphi$ (4bvbwchm)
$=\alpha\varphi=R_{\text{tot}}\varphi^{2}$ (4bvbwchn)
@@
$R_{\text{tot}}$ $=R_{r}+R_{d}$ (4bvbwcho)
$R_{r}$ $=R_{ra}+R_{rb}$ (4bvbwchp)
$R_{d}$ $=R_{d1}+R_{d2}$ (4bvbwchq)
According to Eqs. (4bvbwcc), (4bvbwcd) and (4adc), the total chemical force $\alpha$ can be expressed as:
$\alpha$ $=-\alpha_{ra}+\alpha_{rb}+\alpha_{d1}-\alpha_{d2}$ (4bvbwcia)
$=\ln\left(K_{a}^{-1}\frac{u^{\infty}_{2a}}{u^{\infty}_{1a}}\cdot K_{b}\frac{u^{\infty}_{1b}}{u^{\infty}_{2b}}\cdot\frac{u^{\infty}_{1a}}{u^{\infty}_{1b}}\cdot\frac{u^{\infty}_{2b}}{u^{\infty}_{2a}}\right)$ (4bvbwcib)
$=\ln\left(\frac{K_{b}}{K_{a}}\right)$ (4bvbwcic)
$=\left(\frac{\theta_{g}}{\theta_{g}+1}-\frac{\theta_{g}}{\theta_{g}-1}\right)\eta$ (4bvbwcid)
$=\frac{2\eta\theta_{g}}{1-\theta_{g}^{2}}$ (4bvbwcie)
@@
$\alpha$ $=2\eta\theta_{g}+O(\theta_{g}^{3})\approx 2\eta\theta_{g}$ (4bvbwcj)
@@ C.2 Expression of the steady chemical flux
$\varphi$ $=\frac{\rho\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\frac{\kappa^{\prime}\eta\theta_{g}}{2}$ (4bvbwcka)
@@
$\kappa^{\prime}$ $=\frac{2\kappa}{1+\kappa^{2}}$ (4bvbwckb)
@@
$\lambda^{\prime}$ $=\frac{2\kappa\lambda}{1+\kappa^{2}\lambda^{2}}$ (4bvbwckc)
$\kappa^{\prime}$ and $\lambda^{\prime}$ are parameters that characterize the asymmetry of the chemical reaction and chemical exchange, respectively. The chemical flux in Eq. (4bvbwcka) indicates that the behavior of the system depends on the relative values of $\rho$ and $\lambda^{\prime}\delta$, with the intensity proportional to the reaction asymmetry $\kappa^{\prime}$.
@@
The total entropy production is distributed between the chemical reactions and exchanges:
$\alpha_{r}$ $=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\alpha$ (4bvbwcla)
$\alpha_{d}$ $=\frac{\rho}{\lambda^{\prime}\delta+\rho}\alpha$ (4bvbwclb)
@@
$\sigma_{r}$ $=\frac{\lambda^{\prime}\delta}{\lambda^{\prime}\delta+\rho}\sigma$ (4bvbwcma)
$\sigma_{d}$ $=\frac{\rho}{\lambda^{\prime}\delta+\rho}\sigma$ (4bvbwcmb)
Distribution of the reaction entropy production in each compartment
@@
$\alpha_{ra}$ $=-\frac{1+\varepsilon}{2}\alpha_{r}$ (4bvbwcna)
$\alpha_{rb}$ $=+\frac{1-\varepsilon}{2}\alpha_{r}$ (4bvbwcnb)
@@
$\varepsilon$ $=\beta+\frac{\eta\rho}{2(\lambda^{\prime}\delta+\rho)}\frac{1-\kappa^{2}\lambda^{2}}{1+\kappa^{2}\lambda^{2}}$ (4bvbwcnc)
implying the following distribution of entropy production:
$\sigma_{ra}$ $=\frac{1+\varepsilon\theta_{g}}{2}\sigma_{r}$ (4bvbwcoa)
$\sigma_{rb}$ $=\frac{1-\varepsilon\theta_{g}}{2}\sigma_{r}$ (4bvbwcob)
@@
Distribution of the exchange entropy production for each compound
@@
$\alpha_{d1}$ $=+\frac{\lambda^{2}\kappa^{2}}{1+\lambda^{2}\kappa^{2}}\alpha_{d}$ (4bvbwcpa)
$\alpha_{d2}$ $=-\frac{1}{1+\lambda^{2}\kappa^{2}}\alpha_{d}$ (4bvbwcpb)
implying the following distribution of entropy production:
$\sigma_{d1}$ $=\frac{\lambda^{2}\kappa^{2}}{1+\lambda^{2}\kappa^{2}}\sigma_{d}$ (4bvbwcqa)
$\sigma_{d2}$ $=\frac{1}{1+\lambda^{2}\kappa^{2}}\sigma_{d}$ (4bvbwcqb)